### v3

@@ 1 Introduction
Beyond two dimensions, flip-connectivity is a more challenging concept. For example, the greedy algorithm that flips a locally non-convex configuration succeeds in constructing the Delaunay triangulation in the plane [18], but can get stuck in local minima in three dimensions; see [16] for examples. However, a more limited protocol that inserts points incrementally and repairs the Delaunay triangulation after every insertion also succeeds in three dimensions [17] and extends to higher dimensions and to coherent triangulations [11]. Note that the latter are known in the literature under a variety of names, including Laguerre, regular, and weighted Delaunay triangulations. While flip-connectivity in three dimensions is still an open question, Santos [30] has exhibited a configuration in six dimensions whose family of triangulations is not flip-connected. We refer to [8, 15, 19, 20, 32, 35] and references therein for a multitude of results on flip-connectivity in a variety of settings.
In this paper, we study flips in the family of hypertriangulations of a finite set of $n$ points in the plane. These are triangulations induced by projections of an $(n-1)$-dimensional hypersimplex to the plane. To explain these concepts, we fix an integer $k$ between $1$ and $n-1$, called the level, and we write $\Delta_{n}^{(k)}$ for $k$-fold scaling of the convex hull of the barycenters of the $(k-1)$-dimensional faces of the standard $(n-1)$-simplex, $\Delta_{n}=\Delta_{n}^{(1)}$. Correspondingly, we write ${{A}^{({k})}}$ for the set of $k$-fold sums of the points in the given set of $n$ points, $A={{A}^{({1})}}$. The projection fixes a bijection between the vertices of $\Delta_{n}$ and the points of $A$ and, by construction, maps the vertices of $\Delta_{n}^{(k)}$ to the points of ${{A}^{({k})}}$. The hypertriangulations follow by selecting and projecting appropriate subsets of the $2$-dimensional faces of $\Delta_{n}^{(k)}$. The subclass of coherent level-$k$ hypertriangulations are also known as weighted order-$k$ Delaunay triangulations, defined by generalizing order-$k$ Delaunay triangulations [3] to the weighted setting.
@@
Hypertriangulations were introduced by Olarte and Santos [23] as a tool to study the Baues problem for families of plabic graphs related to the totally positive Grassmannian [26]. The case $k=1$ corresponds to the family of usual triangulations of the set $A$. In the general setting of subdivisions induced by projection between two polytopes, the notion of flips comes from the Baues poset, see [5, 29, 31]. It seems that such flips were first discussed by Billera, Kapranov, and Sturmfels in [5] as possible generalizations for bistellar moves in the setting of the generalized Baues conjecture and related connectivity. We refer to these as combinatorial or Baues flips as it seems difficult to give a geometric description in all possible cases, even for a generic projection. Primarily in connection to the Baues problem, the connectivity of the flip graph for various settings related to the subdivisions induced by projections of polytopes has been studied in [2, 5, 21, 28], where positive as well as negative results are described. We also note that a particular family of Baues flips, the $\pi$-flips, was studied by Santos [31].
@@
Noting that we restrict ourselves to generic point sets in the plane, here is the outline of this paper: Section 2 follows Olarte and Santos [23] and introduces level-$k$ hypertriangulations of a set of $n$ points, $A$, as tight subdivisions induced by the projection of the hypersimplex $\Delta_{n}^{(k)}$. In particular, we give a combinatorial description of such triangulations without using the associated projection. We also define the more general hypersimplicial subdivisions and sketch a connection to fiber polytopes [6]. Section 3 introduces the four types of flips for level-$k$ hypertriangulations. Notably, these flips are geometric, they do not depend on the level $k$, and they include the classic flips for triangulations of planar point sets [7].
Section 4 proves the main result of this paper: that the family of level-2 hypertriangulations of every (generic) planar point set is flip-connected. The main tool in the proof is the aging function for triangles of hypertriangulations defined in Section 4.1. We note that the aging function was used before by Galashin [12] and Olarte and Santos [23] for point sets in convex position. For our approach, properties of the aging function for arbitrary generic point sets are instrumental and we prove these properties in Sections 4.1 and 4.2.
Section 5 discusses connections of flips in hypertriangulations to plabic graphs studied by Postnikov [25, 26] and Galashin [12], and reformulates their results in the language of flips. It also discusses flips within the family of coherent triangulations and the relation to the fiber polytopes, and the connection of flips we introduce to $\pi$-flips by Santos [31]. Section 6 concludes the paper with a discussion of possible further research.

### v4

@@ 1 Introduction
Beyond two dimensions, the connectivity with flips also called Pachner moves or bistellar flips, is a more challenging concept. For example, the greedy algorithm that flips a locally non-convex configuration succeeds in constructing the Delaunay triangulation in the plane [18], but can get stuck in local minima in three dimensions; see [16] for examples. However, a more limited protocol that inserts points incrementally and repairs the Delaunay triangulation after every insertion also succeeds in three dimensions [17] and extends to higher dimensions and to coherent triangulations [11]. Note that the latter are known in the literature under a variety of names, including Laguerre, regular, and weighted Delaunay triangulations. While flip-connectivity in three dimensions is still an open question, Santos [30] has exhibited a configuration in six dimensions whose family of triangulations is not flip-connected. We refer to [8, 15, 19, 20, 32, 35] and references therein for a multitude of results on flip-connectivity in a variety of settings.
@@
In a more general setting beyond triangulations, flips were introduced in the study of subdivisions induced by projections between two polytopes and associated fiber polytopes by Billera and Sturmfels [6]. The corresponding notion of flips comes from the Baues poset, see [5, 29, 31]. It seems that such flips were first discussed by Billera, Kapranov, and Sturmfels in [5] as possible generalizations for bistellar moves in the setting of the generalized Baues conjecture and related connectivity. We refer to these as combinatorial or Baues flips as it seems difficult to give a geometric description in all possible cases, even for a generic projection. Primarily in connection to the Baues problem, the connectivity of the flip graph for various settings related to the subdivisions induced by projections of polytopes has been studied in [2, 5, 21, 28], where positive as well as negative results are described.
In this paper, we study flips in the family of hypertriangulations of a finite set of $n$ points in the plane. Hypertriangulations were introduced by Olarte and Santos [23] as a tool to study the Baues problem for families of plabic graphs related to the totally positive Grassmannian [26]. These are triangulations induced by projections of an $(n-1)$-dimensional hypersimplex to the plane. To explain these concepts, we fix an integer $k$ between $1$ and $n-1$, called the level, and we write $\Delta_{n}^{(k)}$ for the $k$-fold scaling of the convex hull of the barycenters of the $(k-1)$-dimensional faces of the standard $(n-1)$-simplex, $\Delta_{n}=\Delta_{n}^{(1)}$. Correspondingly, we write ${{A}^{({k})}}$ for the set of $k$-fold sums of the points in a given set of $n$ points, $A={{A}^{({1})}}$. The projection fixes a bijection between the vertices of $\Delta_{n}$ and the points of $A$ and, by construction, maps the vertices of $\Delta_{n}^{(k)}$ to the points of ${{A}^{({k})}}$. The hypertriangulations follow by selecting and projecting appropriate subsets of the $2$-dimensional faces of $\Delta_{n}^{(k)}$. The subclass of coherent level-$k$ hypertriangulations are also known as weighted order-$k$ Delaunay triangulations, defined by generalizing order-$k$ Delaunay triangulations [3] to the weighted setting. The case $k=1$ corresponds to the family of usual triangulations of the set $A$ that can be viewed as (tight) subdivisions induced by the projection of (the vertices of) the standard simplex $\Delta_{n}$ on the points of $A$. Our main result establishes flip-connectivity of level-2 hypertriangulations of generic point sets.
Theorem 4.4.
The flips we consider in the theorem above are possibly more restrictive than the Baues flips. They are introduced geometrically in Section 3. For the specific setting of projections of hypersimplices, these flips coincide with $\pi$-flips introduced by Santos in [31] as we briefly discuss in Section 5.1 and are subsets of the Baues flips for the corresponding projections. However, we were unable to prove or disprove that the family of flips we use coincide with the family of Baues flips for (generic) projections of hypersimplices.
Since hypersimplices are simplicial polytopes, the family of level-$k$ hypertriangulations of $A$ is a subset of the family of triangulations of the set $A^{(k)}$. Notwithstanding, the flip connectivity of the former family does not follow from the flip-connectivity of the latter family and the flips are defined differently. Moreover, flip-connectivity is not guaranteed for projections of simplicial polytopes on the plane as illustrated by the example of Rambau and Ziegler [28]. Thus, in this paper we prove that every (generic) projection of the hypersimplex $\Delta_{n}^{(2)}$ gives rise to a connected Baues poset extending the same property known for projections of simplices.
@@
Before presenting the outline of this paper, we stress that we restrict ourselves to generic point sets in the plane. Section 2 follows Olarte and Santos [23] and introduces level-$k$ hypertriangulations of a set of $n$ points, $A$, as tight subdivisions induced by the projection of the hypersimplex $\Delta_{n}^{(k)}$. In particular, we give a combinatorial description of such triangulations without using the associated projection. We also define the more general hypersimplicial subdivisions and sketch a connection to fiber polytopes [6]. Section 3 introduces the four types of flips for level-$k$ hypertriangulations. Notably, these flips are geometric, they do not depend on the level $k$, and they include (colored versions of) the classic flips for triangulations of planar point sets [7]. Section 4 proves the main result of this paper: that the family of level-2 hypertriangulations of every (generic) planar point set is flip-connected. The main tool in the proof is the aging function for triangles of hypertriangulations defined in Section 4.1. We note that the aging function was used before by Galashin [12] and Olarte and Santos [23] for point sets in convex position. For our approach, properties of the aging function for arbitrary generic point sets are instrumental and we prove these properties in Sections 4.1 and 4.2. Section 5 discusses relations between flips in hypertriangulations to plabic graphs studied by Postnikov [25, 26] and Galashin [12], and reformulates their results in the language of flips. It also discusses flips within the family of coherent triangulations and the relation to the fiber polytopes, and the connection of flips we introduce to $\pi$-flips by Santos [31]. Section 6 concludes the paper with a discussion of possible further research.